An object has a rectangular shape with a surface area of 40m2 and 0.1m thick. What is the maximum mass of a polar bear can sit on the floe before sinking? (density of ice is 900kg/m3 and density of seawater is 1025kg/m3)


1
Expert's answer
2021-01-29T16:57:04-0500

"S=40~m^2,~h=0.1~m,~\rho_i=900~\frac{kg}{m^3},~\rho_w=1025~\frac{kg}{m^3}."


"P=m_ig=\rho_igSh,"

"F_A=\rho_wgV'=\rho_wgSh',"

"P=F_A,"

"\rho_ih=\rho_wh',"

"h'=h\frac{\rho_i}{\rho_w},"

"m_{bear}=m_{upper~ice}=\rho_iS(h-h')=\rho_iSh(1-\frac{\rho_i}{\rho_w}),"

"m_{bear}=900\cdot 40\cdot 0.1\cdot(1-\frac{900}{1025})=439 ~kg."